### 03 Mar 2024 v1

# Fine Tuning vs. Retrieval Augmented Generation for Less Popular Knowledge

Abstract
Large language models (LLMs) memorize a vast amount of factual knowledge, exhibiting strong performance across diverse tasks and domains. However, it has been observed that the performance diminishes when dealing with less-popular or low-frequency concepts and entities, for example in domain specific applications. The two prominent approaches to enhance the performance of LLMs on low-frequent topics are: Retrieval Augmented Generation (RAG) and fine-tuning (FT) over synthetic data. This paper explores and evaluates the impact of RAG and FT on customizing LLMs in handling low-frequency entities on question answering task. Our findings indicate that FT significantly boosts the performance across entities of varying popularity, especially in the most and least popular groups, while RAG surpasses other methods. Additionally, the success of both RAG and FT approaches is amplified by advancements in retrieval and data augmentation techniques. We release our data and code at https://github.com/HeydarSoudani/RAGvsFT.
Fine Tuning vs. Retrieval Augmented Generation for Less Popular Knowledge
Heydar Soudani Radboud University Nijmegen The Netherlands heydar.soudani@ru.nl                        Evangelos Kanoulas University of Amsterdam Amsterdam The Netherlands e.kanoulas@uva.nl                        Faegheh Hasibi Radboud University Nijmegen The Netherlands f.hasibi@cs.ru.nl
1 Introduction
Large Language Models (LLMs) exhibit outstanding capabilities in executing tasks that demand extensive memorization of factual data (Chowdhery et al., 2023). However, their memorization capabilities are constrained when dealing with less frequent entities Mallen et al. (2023); Kandpal et al. (2023); Sun et al. (2023), and even the largest models may encounter the well-known "hallucination" problem Shuster et al. (2021) and temporal degradation Kasai et al. (2022). Consequently, when LLMs are intended for deployment in less resourced domains, customization becomes imperative to ensure optimal performance. A common example is within the industrial setup, where chatbots or Question Answering (QA) systems need to accurately answer users’ questions about a proprietary knowledge graph or intra-company terminology with limited textual description.
Retrieval-Augmented Generation (RAG) and Fine-Tuning (FT) stand out as two prominent approaches for adapting LLMs to specific domains. RAG retrieves relevant information from a document corpus and enhances LLM’s response generation through the implementation of in-context learning (ICL). Conversely, the fine-tuning approach updates model weights to become adept at recalling specific information and enhance its memorization capabilities during inference. In the context of less popular knowledge, where limited data is available, data augmentation methods are utilized to generate synthetic training data, serving as an initial step towards fine tuning.
Figure 1: Correlation between subject entity popularity in a question and the effects of RAG and FT on FlanT5-small performance in open-domain question answering. FT markedly improves accuracy in the initial and final buckets relative to others (indicated by the pink line).
In this paper, we aim to understand which approach and under what conditions is more effective for industry-specific models. Specifically, we seek to answer the following research questions:
(RQ1) What is the effectiveness of RAG and fine-tuning with synthetic data on QA for low-frequency factual knowledge?
(RQ2) Which parameters, including the quality of synthetic samples, the method to be fine-tuned, the model size, and the performance of retrieval models affect the downstream performance?
To address our RQs, we performed a comprehensive comparison of RAG and fine tuning methods, with specific attention to less popular knowledge. Our evaluation setup explores various factors, including model size, retrieval models, the quality of synthetic data generation, and the fine-tuning method (PEFT vs. full fine tuning).
Our findings indicate that FT consistently enhances performance for entities, both popular and less popular, with the most substantial improvements observed in the most and least popular categories. Furthermore, RAG consistently outperforms fine tuning methods, particularly when combined with FT in smaller models, a benefit that diminishes in base models and is non-existent in larger models. Lastly, the effectiveness of both RAG and FT strategies increases with improvements in the performance of the retrieval and data augmentation models.
2 Background
Data Augmentation (DA). Data availability is crucial for fine-tuning in specialized domains. DA addresses data scarcity problem by generating task- and domain-relevant samples from existing unlabeled texts. A common DA approach for the QA task is generating question-answer pairs through a four-step Pipeline, consisting of: passage selection, answer extraction, question generation, and consistency filtering (Alberti et al., 2019; Lewis et al., 2019, 2021; Ushio et al., 2023). Ushio et al. (2023) conducted an empirical study comparing three question answer generation approaches: Pipeline, Multitask, and End-to-End (E2E) and showed the E2E approach outperforms others in downstream tasks. Recently, the utilization of LLMs to generate data is shown effective in information retrieval, QA, and dialogue creation tasks (Soudani et al., 2023; Askari et al., 2023).
Retrieval Augmented Generation. RAG enhances LLMs by integrating external knowledge sources with input queries, enriching the model with additional context for knowledge-intensive tasks (Lewis et al., 2020). It utilizes an information retrieval system to find relevant documents and adds them to the input prompt to enhance response generation of an LLM (Asai et al., 2023a, b).
Less popular Knowledge. An entity’s popularity in an LLM is gauged by its frequency in the model’s pre-training data (Godbole and Jia, 2023; Min et al., 2023), often assessed through the entity’s occurrence in a large corpus (Kandpal et al., 2023). Due to the practical challenges of direct counting, approximations like traffic metrics and content density are used (Sun et al., 2023). Wikipedia pageviews are among the most prevalent methods for measuring the popularity of entities (Mallen et al., 2023; Sciavolino et al., 2021; Chen et al., 2021).
Comparing FT vs. RAG. As interest grows in refining pre-trained language models for particular tasks, the comparison of FT and RAG strategies under equitable conditions is becoming increasingly important. Mosbach et al. (2023) explored the effectiveness of few-shot FT versus In-context Learning for classification tasks in general domains. de Luis Balaguer et al. (2024) compared FT and RAG in answering long, agriculture, and geography-specific questions. Ovadia et al. (2023) assessed the performance on multiple-choice questions in specialized areas like Anatomy, Astronomy, College Biology, and Prehistory. In contrast to these studies, we directly address the integration of less popular factual knowledge into LLMs, comparing various retrieval, data augmentation, and fine tuning methods.
3 Evaluation Setup
We evaluate LLMs on the closed-book QA task, focusing on the PopQA dataset (Mallen et al., 2023) characterized by questions that covers a long-tail entity distribution. It also includes unique Wikipedia titles, facilitating the calculation of pageviews and identification of relevant Wikipedia pages. We acquire the relevant Wikipedia page for each subject entity in the PopQA dataset and divide the entities into five buckets based on their popularity levels (Figure 2).
Figure 2: Distribution of sample counts across buckets, defined by $log_{10}(\text{pageviews})$. The leftmost bin includes entities with fewer than $10^{2}$ pageviews, while the rightmost bin encompasses entities with over $10^{5}$ pageviews.
To ensure a fair comparison between DA and RAG methods, we limit our focus to Wikipedia pages whose corresponding entities appear in the PopQA dataset. This setup also mirrors real-world industry practices, where entities and their corresponding textual descriptions are related to companies’ specific internal concepts.
3.1 RAG Approach
We utilize a variety of retrieval models to retrieve relevant passages for the RAG approach: BM25 (Robertson and Zaragoza, 2009), Contriever (Izacard et al., 2022), DPR (Karpukhin et al., 2020), and a two-stages re-ranker that combines BM25 with DPR, all implemented according to the BEIR benchmark (Thakur et al., 2021). Moreover, we utilize an ideal retrieval setup, where the summary paragraph of the Wikipedia page for the mentioned entity in the question is considered. We adopt this method because the true evidence for annotating passages is not available, yet the summary section assumably contains the answer to the question. The efficacy of this approach is evidenced in Figure 5 in Appendix D, which shows QA performance with the the ideal retriever achieves the highest results.
After retrieving the most relevant passage for a question, we apply zero-shot prompting for generative prediction using a straightforward template: "Context: <context>. Based on the provided
context, answer the question: <question>". We experiment with three sizes of the FlanT5 model (Chung et al., 2022)—small, base, and large—to examine the impact of model scale on performance. Following Mallen et al. (2023), we report on the Accuracy metric, where a prediction is considered accurate if it contains a substring that exactly matches any of the provided gold answers.
3.2 Fine-Tuning Approach
We generate training data for the fine tuning approach using two distinct data augmentation methods. The first one is the End-to-End approach (Ushio et al., 2023), utilizing a model specifically trained for paragraph-level QA generation, using the T5 (Raffel et al., 2020) model, referred to as E2E in our paper. Additionally, we explore the generation of synthetic training data by prompting an LLM, utilizing Zephyr (Tunstall et al., 2023) for QA generation. This method is referred to as the Prompt method in our discussion. It should be noted that we generate QAs exclusively using the summary section of Wikipedia pages. This approach ensures a fair comparison between FT and RAG with ideal retrieval. Further details on the QA generation process, including the prompt, input and output examples, and statistics of the generated QAs, are provided in the Appendix B.
After generating QA pairs, we proceeded to fine-tune the FlanT5 models using two approaches: full parameter tuning (Full) and Parameter Efficient Fine-Tuning (PEFT). Within the range of PEFT techniques (Zaken et al., 2022; Liu et al., 2022; Ma et al., 2023), we utilize QLoRA (Dettmers et al., 2023), chosen for its broad acceptance in the field (Kaddour et al., 2023; Naveed et al., 2023) and efficient use of computational resources we had at our disposal. More information about FT approaches and hyper parameters are detailed in Appendix C.
4 Experiments and Results
Model Rec@1 Rec@3 Rec@5
BM25 36.14 73.76 81.68
Contriever 72.77 91.58 96.04
BM25+DPR 79.21 91.58 92.57
DPR 78.21 92.57 93.64
Table 1: Performance comparison of retrieval models on our corpus, with DPR models outperforming the rest. The relevant passage for every question is assumed to be the first paragraph of the Wikipedia page.
Retrieval performance. Figure 3 compares the performance of retrieval models against ideal retriever (assumed to be the ground truth here) across popularity buckets. Since only the highest-ranked passage is included in the input prompt, our analysis is focused on Recall@1. The results indicate that retrieval effectiveness is higher in low popular entities compare to popular entities. This is probably due to limited occurrences of noisy passages for low popular entities. Table 1 provides the overall retrieval scores for all methods.
FT and RAG comparison. We assess the impact of RAG and FT in four distinct configurations: (i) Neither FT nor RAG used (-FT-RAG), (ii) RAG used without FT (-FT+RAG), (iii) FT applied without RAG (+FT-RAG), and (iv)) both FT and RAG employed (+FT+RAG). The findings are detailed in Table 2. Generally, FT improves accuracy of the base model, but does not reach the effectiveness of RAG on the base model. The optimal performance is achieved by integrating both FT and RAG.
-FT/ -FT/ +FT/ +FT/
 FT QA -RAG +RAG -RAG +RAG
FlanT5-small
 PEFT E2E 3.05 26.13 5.53 22.91
Prompt 7.01 49.85
 Full E2E 6.35 10.21
Prompt 8.52 49.88
FlanT5-base
 PEFT E2E 6.72 63.13 6.94 51.61
Prompt 9.92 63.29
 Full E2E 8.63 24.17
Prompt 11.41 60.26
FlanT5-large
 PEFT E2E 8.41 58.12 8.22 55.26
Prompt 12.15 61.71
 Full E2E 16.23 13.31
Prompt 13.91 58.60
Table 2: Accuracy of base and fine-tuned models, both with and without RAG, is reported across various conditions, encompassing model size, fine-tuning approaches, and QA generation techniques. The RAG results presented are based on ideal retrieval.
Figure 3: Recall@1 performance of four retrieval methods across different popularity levels of pageviews. The results indicate that retrieval models perform more effectively with less popular knowledge compared to more popular ones.
Figure 1 demonstrates that fine-tuning enhances QA accuracy across all popularity levels for the FlanT5-small model, with the greatest improvements in the most and least popular buckets. While fine-tuning boosts performance for less popular categories in smaller models, this advantage decreases in base models and disappears in larger models; see Figure 4 for more details. This is likely due to the improved memorization of larger models.
Effect of external parameters on RAG and FT. We delve into additional factors that influence model specialization in processing less popular knowledge. A key aspect under review is the effect of full tuning versus PEFT. Table 2 shows PEFT leads to smaller gains in the +FT-RAG compared to full FT, yet it significantly improves accuracy in the +FT+RAG setup. This suggests that PEFT enables the LLM to maintain its reasoning abilities based on the provided prompts.
Our investigation also covers two QA generation techniques. The E2E method generates over 12 times more QA than the prompting method (cf. Table 4 in Appendix C), while prompt-based method generates more quality QAs. The results in Table 2 show that fine-tuned models trained on prompt-generated data outperform E2E-generated ones. This highlights the significance of synthetic data quality over quantity.
The performance of the retrieval model is another important consideration. Table 3 presents the QA system’s accuracy utilizing various retrieval strategies within the RAG framework. The findings demonstrate a direct correlation between the retrieval model performance and the overall QA accuracy, underscoring the retrieval model’s impact on the downstream task’s effectiveness.
 FT QA
BM25
Contriever
BM25+DPR
DPR
Ideal
FlanT5-small 9.76 17.22 20.80 21.41 26.13
 PEFT E2E 13.13 17.30 19.74 19.81 22.91
Prompt 20.82 34.87 40.45 41.25 49.85
 Full E2E 6.48 9.11 9.31 9.48 10.21
Prompt 20.95 35.09 40.52 41.53 49.88
FlanT5-base 26.56 43.42 50.41 51.35 63.13
 PEFT E2E 22.72 35.80 42.05 42.84 51.61
Prompt 26.83 43.20 50.59 51.43 63.29
 Full E2E 11.78 18.53 20.59 21.15 24.17
Prompt 25.77 41.78 48.52 49.39 60.26
FlanT5-large 26.74 40.53 46.86 47.57 58.12
 PEFT E2E 25.72 39.25 45.34 45.90 55.26
Prompt 28.60 43.45 50.15 50.91 61.71
 Full E2E 7.67 12.29 12.01 12.44 13.31
Prompt 25.54 40.88 46.82 47.87 58.60
Table 3: Accuracy of RAG for base and fine-tuned LLMs using different retrieval models.
5 Conclusion
In this study, we performed a comparative analysis to evaluate the effectiveness of RAG versus FT, with a focus on less popular knowledge. Our results reveal that FT leads to consistent performance improvements for all entities, with the most notable gains seen in the most and least popular categories. We found that RAG stands out as a more effective strategy, especially when used in combination with fine-tuning. This advantage decreases however in larger models. Additionally, we observed that the success of both RAG and FT strategies improves with the enhancement of the retrieval and data augmentation models’ performance. Understanding the critical importance of synthetic data quality, future work will focus on developing an effective method for data creation.
6 Limitations
Our study primarily addressed a template-based QA task, suggesting future research could tackle more complex QA challenges, such as multi-hop QA (Ho et al., 2020) or Conversational QA (Christmann et al., 2022). Furthermore, we noted that our use of Zephyr for synthetic data generation has limitations in following Chain of Thought (CoT), highlighting the potential benefits of advanced data generation techniques to enhance data quality and reduce fine-tuning cost.
Appendix
Appendix A Data Preprocessing
Among the various datasets offering factual QA (Christmann et al., 2023; Sciavolino et al., 2021), we selected one of the most recent ones, named PopQA (Mallen et al., 2023). PopQA dataset comprises approximately 14,000 templated questions aimed at eliciting single-entity answers, based on 16 types of relationships extracted from Wikipedia. To assess the popularity of each entity, we followed the previous studies (Mallen et al., 2023; Sciavolino et al., 2021; Chen et al., 2021), using Wikipedia pageviews as a metric for popularity. We accumulated the pageviews for each entity from the start to the end of 2023.
Furthermore, we acquired the relevant Wikipedia page for each entity. Our corpus was then constructed by segmenting each Wikipedia page into a summary paragraph and subsequent additional paragraphs. For this purpose, we utilized the Wikipedia dump from March 2022, available in the HuggingFace dataset repository at https://huggingface.co/datasets/wikipedia.
Appendix B QA Generation
B.1 Prompt Method
For generating synthetic QAs based on a given context, we used the following prompt: You are a question-answer generator. Your goal is to generate question-answer pairs given the Context.
Example output: {’question’: ”, ’answer’: ”}
Context: \$CONTEXT
Step 1: Identify spans that are likely to be answers to questions, identify as many as possible. Step 2: For each identified span, generate a question. Step 3: Respond to the question in only a few tokens concisely. Step 4: Output in JSON format following […]
Ensure that you distinctly label and delineate Steps 1, 2, 3, and 4. Let’s think step by step:
Our test set comprises template-based QAs that feature single-entity answers. To generate QAs similar to those in the test set, we instructed the LLM to utilize a Chain of Thought (CoT) approach. However, we observed that not all the generated QAs adhered to our instructions. Below are some examples of QAs generated using this prompt-based approach:
Entity: John Mayne
Context: John Mayne was a Scottish printer, journalist and poet born in Dumfries. In 1780, his poem The Siller Gun appeared in its original form in Ruddiman’s Magazine, published by Walter Ruddiman in Edinburgh. It is a humorous work on an ancient custom in Dumfries of shooting for the S̈iller Gun.Ḧe also wrote a poem on Hallowe’en in 1780 which influenced Robert Burns’s 1785 poem Halloween. Mayne also wrote a version of the ballad Helen of Kirkconnel. His verses were admired by Walter Scott.
Synthetic QAs: question 1: "Who was John Mayne?" answer 1: "John Mayne was a Scottish printer, journalist, and poet from Dumfries. (1780 poem: The Siller Gun)"
question 2: "What poem did John Mayne write in 1780 called Hallowe’en?", answer 2: "Hallowe’en is a poem by John Mayne (1780)"
B.2 E2E Method
We employed models trained for the Question Answer Generation (QAG) task, specifically using the T5-large model optimized for the End-to-End (E2E) approach, which was reported to deliver superior performance on the QA as the downstream task (Ushio et al., 2023). However, a notable drawback of the E2E method is its limited control over the quality and quantity of generated samples. Additionally, due to the T5 model’s input token limitation, the input text must be restricted to 512 tokens. To ensure a fair comparison, we applied the same token limit to the inputs used with the prompt-based approach. Table 4 presents statistics for the QAs generated by both the prompt and E2E methods, using the same corpus.
Relationship #Ent. Train-set Dev-set
E2E Prompt E2E Prompt
occupation 532 16398 1872 1822 207
place of birth 584 19993 2043 2221 227
genre 1619 56101 5338 6233 593
father 570 37676 2092 4186 406
country 838 19362 2769 2151 307
producer 1520 41992 4235 4665 470
director 1999 27578 4644 3064 515
capital of 363 109141 3388 12126 376
screenwriter 1999 59680 6932 6631 770
composer 978 46574 4064 5174 451
color 34 4396 160 488 17
religion 338 18776 1449 2086 161
sport 547 14760 1710 1629 189
author 1514 46399 5319 5155 590
mother 187 7592 477 843 52
capital 645 28467 1322 3162 146
All 14,267 491,525 38,114 56,803 5,949
Table 4: Statistics for generated QA pairs using two methods. The E2E approach, utilizing T5-large, generates over 12 times more QAs compared to the prompt method.
Appendix C Fine-Tuning Setup
We fine-tuned three versions of FlanT5 (small, base, and large) (Chung et al., 2022) model. Our fine-tuning experiments encompassed both full parameter tuning and PEFT. Our focus was particularly on the QLoRA method (Dettmers et al., 2023), which is widely recognized for its approach of keeping the pre-trained model parameters fixed while incorporating trainable rank decomposition matrices into each layer of the Transformer architecture. Despite experimenting with various numbers of training epochs, we observed that the results plateaued after a certain point. Table 5 included presents our finalized hyperparameters for fine-tuning.
Hyperparameter Value
Epochs 10
Batch size 128
Learing rate 2e-4
PEFT, alpha 16
PEFT, rank 32
PEFT, dropout 0.05
Table 5: Hyperparameters for Fine-Tuning: Following a thorough hyperparameter search, we standardized the hyperparameters across all versions of the model.
Appendix D Detailed Results
In this section, we provide additional results that corroborate our primary findings. Figure 4 illustrates the impact of fine-tuning versus non-fine-tuning, both with and without RAG, on the FlanT5-base and FlanT5-large models. It further demonstrates that fine-tuning enhances accuracy across both the least-popular and popular buckets. It should be noted that for the base model, the improvement in accuracy for the least-popular bucket is less significant compared to that of the small and large models.
Figure 4: Correlation between the popularity of the subject entity in a question and the impact of RAG and FT on the performance of FlanT5-base and FlanT5-large in QA.
Figure 5 displays the accuracy of the RAG approach before fine-tuning across various retrieval models. The Ideal retriever, which retrieves the summary section of the corresponding Wikipedia page, yields the highest accuracy.
Figure 5: The accuracy of the RAG approach before fine-tuning across all relationships in PopQA dataset.
Figure 6 compares the performance before and after fine-tuning, both with and without RAG, across all relationships. It also demonstrates that fine-tuning alone does not achieve the same level of accuracy as the RAG method for most of relationships.
Figure 6: The accuracy of the RAG approach before and after fine-tuning across all relationships in PopQA dataset.

### 07 Mar 2024 v2

@@
 Additionally, the success of both RAG and FT approaches is amplified by advancements in retrieval and data augmentation techniques.
-We release our data and code at https://github.com/HeydarSoudani/RAGvsFT.
+The code and data is available at https://github.com/informagi/RAGvsFT.
 Fine Tuning vs. Retrieval Augmented Generation for Less Popular Knowledge
-Heydar Soudani Radboud University Nijmegen The Netherlands heydar.soudani@ru.nl                        Evangelos Kanoulas University of Amsterdam Amsterdam The Netherlands e.kanoulas@uva.nl                        Faegheh Hasibi Radboud University Nijmegen The Netherlands f.hasibi@cs.ru.nl
+Heydar Soudani Radboud University Nijmegen The Netherlands heydar.soudani@ru.nl                        Evangelos Kanoulas University of Amsterdam Amsterdam The Netherlands e.kanoulas@uva.nl                        Faegheh Hasibi Radboud University Nijmegen The Netherlands faegheh.hasibi@ru.nl
 1 Introduction
@@
 We generate training data for the fine tuning approach using two distinct data augmentation methods.
-The first one is the End-to-End approach (Ushio et al., 2023), utilizing a model specifically trained for paragraph-level QA generation, using the T5 (Raffel et al., 2020) model, referred to as E2E in our paper.
+The first one is the End-to-End approach (Ushio et al., 2023), utilizing a model specifically trained for paragraph-level QA generation, using the T5 model (Raffel et al., 2020), referred to as E2E in our paper.
 Additionally, we explore the generation of synthetic training data by prompting an LLM, utilizing Zephyr (Tunstall et al., 2023) for QA generation.
 This method is referred to as the Prompt method in our discussion.
-It should be noted that we generate QAs exclusively using the summary section of Wikipedia pages.
-This approach ensures a fair comparison between FT and RAG with ideal retrieval.
+We generate QAs exclusively using the summary section of Wikipedia pages to ensures a fair comparison between FT and RAG with ideal retriever.
 Further details on the QA generation process, including the prompt, input and output examples, and statistics of the generated QAs, are provided in the Appendix B.
@@
 Prompt 13.91 58.60
-Table 2: Accuracy of base and fine-tuned models, both with and without RAG, is reported across various conditions, encompassing model size, fine-tuning approaches, and QA generation techniques.
-The RAG results presented are based on ideal retrieval.
-Figure 3: Recall@1 performance of four retrieval methods across different popularity levels of pageviews.
+Table 2: Accuracy of base and fine-tuned models, both with and without RAG. The RAG results presented are based on ideal retrieval.
+Figure 3: Recall@1 for four retrieval models across different popularity levels.
 The results indicate that retrieval models perform more effectively with less popular knowledge compared to more popular ones.
@@
 Furthermore, we noted that our use of Zephyr for synthetic data generation has limitations in following Chain of Thought (CoT), highlighting the potential benefits of advanced data generation techniques to enhance data quality and reduce fine-tuning cost.
+Acknowledgements
+This publication is part of the project LESSEN with project number NWA.1389.20.183 of the research program NWA ORC 2020/21 which is (partly) financed by the Dutch Research Council (NWO).
 Appendix